OpenStudy (anonymous):

Assume 3 in every 1000 students at a local community college have quit due to serious health issues. An insurance company offers them an $8000 policy for $50 a year. What is the amount the insurance company should expect to make on average on every student that pays?

OpenStudy (slaaibak):

Income per year: 1000 * 50 = 50 000 Expected claims/payments: 8000 * 3 = 24 000 Expected profit: 50 000 - 24 000 = 26 000